# Vendi Information Gain: An Alternative To Mutual Information For Science And Machine Learning

### v1

@@ 1 Introduction
In Section 2, we demonstrate that VIG overcomes MI’s limitations and outperforms it in diverse settings. We also introduce in Section 2 a novel VIG-based active data acquisition framework, delivering the first fully information-theoretic method for level-set estimation. Finally, we derive VIG’s key properties in Section 5.
2 Results
In this section, we define VIG and then demonstrate its superiority over MI in several applications. In particular, we use VIG to develop a unified framework for active data acquisition, including a novel, first-ever information-theoretic approach to level-set estimation.
2.1 Defining the Vendi information gain
@@
2.2 Vendi information gain addresses mutual information’s failures
@@ 2.3 Modeling human response times to external stimuli
Figure 2: Top: Progression of posterior probabilities and information gain, as measured by MI and VIG, under varying degrees of inter-message similarity as a function of time. MI plots the same trend across different conditions, while VIG becomes lower under more similar messages, appropriately quantifying increasing difficulty in distinguishing similar messages. Bottom: Simulated response times as a function of similarity between possible messages by MI and VIG against observed experimental data from Slamecka (1963). MI’s insensitivity to inter-message similarity prevents the simulation of different behaviors under different conditions, while VIG’s simulations closely match observed real-world data.
@@
We demonstrate VIG’s ability to model the concept of confusability by performing the simulation with the model by Christie et al. (2023), transmitting three symbols $A$, $B$, and $C$ with varying degrees of similarity. In particular, symbol $C$ is modeled as being distinct from $A$ and $B$, while the similarity between $A$ and $B$ varies from low to intermediate to high. Symbol $A$ is set to be the intended stimulus, and the top panel of section 1 shows the posterior probability that each symbol is the intended stimulus $p(m\mid D_{t})$ as the function of time $t$. Here, we start out with a flat prior where each symbol has a $1/3$ probability of being the intended stimulus, and as we observe more data as time goes on, $p(m=A\mid D_{t})$ increases, indicating that $m=A$ is the true message.
The remaining panels of the top plot of section 1 show the progression of IG as measured by MI and VIG in each of the three aforementioned settings. MI treats individual stimuli as completely distinct from one another and plots the same trend across the settings, failing to distinguish the different conditions. Meanwhile, VIG appropriately quantifies increasing difficulty in distinguishing messages as they become more similar, as demonstrated by lower IG. (In the extreme case of low similarity where there is no inter-stimulus similarity, VIG reduces to MI.)
Finally, we repeat the same simulation 100 times and show in the bottom plot of section 1 the distribution of simulated response times resulting from using MI and VIG, as well as the real-world data reported by Slamecka (1963) from their word-meaning experiments as a reference. From the real-world experimental data, we observe the natural trend that the higher the similarity between the stimuli, the harder it is to distinguish between them, and thus the longer the response time is. This trend cannot be replicated using MI but is faithfully simulated by VIG, demonstrating VIG’s advantage.
@@ 2.4 Vendi information gain leads to a general unified framework for active data acquisition
Finally, we perform ablation studies on the effect of the order $q$ has on VIG in figs. 8 and 9 and find that performance remains fairly robust across a wide range of values for $q$, suggesting that VIG’s main advantage over other policies stems from its IG reasoning for active data acquisition tasks.
@@ 3 Related Work
Mutual information.  Previous works have investigated the failure modes of MI. Indeed, researchers have highlighted the challenge in accurately estimating MI in an unbiased way, especially in high dimensions (Macke et al., 2011; Dorval, 2011; Rhee et al., 2012; Kinney and Atwal, 2014; Cheng and Vlachos, 2024). Many studies have also found MI estimators to take on undesirable pathological behaviors under multimodal, non-Gaussian distributions (Czyż et al., 2023; Peyrard and Cho, 2025), and even the accurate identification of these undesirable behaviors of MI is still an active area of search. Approximating MI may require exponentially many samples (Quian Quiroga and Panzeri, 2009), and when it can be approximated with high fidelity or computed exactly, MI’s insensitivity to random variables’ spread may lead to a mischaracterization of IG (Petty, 2018; Oliver, 2022; Cheng and Vlachos, 2024). We echo the findings by Schroeder (2004), who argued that MI is a good IG metric for communication systems when there is no need to model similarity between transmitted symbols, but fails in other contexts.
@@
Various works have considered alternatives to MI that address its drawbacks. However, most of these proposals are tailored to specific applications, and thus do not generalize to a variety of tasks where MI is employed. For instance, Torkkola (2003) considered a version of MI that accounts for the similarity between pairs of samples (something VIG inherently does) using a Parzen estimator as part of a feature extraction procedure. Similarly, for the task of image registration, Haber and Modersitzki (2005) proposed modifying MI to take as input a distance measure and calculated a distance-aware MI-based quantity between two images; the authors showed that this alternative outperforms MI under multi-modal data. In hierarchical clustering where MI has been found to inappropriately favor large clusters under high dimensions, Marrelec et al. (2015) studied an MI alternative that uses the empirical covariance matrix, like VIG, to correct the undesirable cluster-favoring behavior, and ultimately leads to better performance, while Hafemann et al. (2024) addressed MI’s insensitivity to spread by incorporating each data point’s distances to its nearest neighbors when analyzing LiDAR sensors. Meister et al. (2024) computed a similarity-weighted version of Shannon entropy to model contextual word predictability. Similar to findings in our response time application, modeling similarity between words better bridges the gap between information theoretic simulations and real-world data from cognitive science compared to MI. Liu et al. (2022) employed a distance-aware measure that also uses the covariance matrix of the samples to sidestep MI’s intractability for learning fair representations. Kala (2021) pointed out that differential entropy may take on negative values, resulting in unintuitive IG values that are undesirable for sensitivity analysis tasks when quantified by MI, and Petty (2018) made similar observations in the context of modeling atmospheric variables. Vendi entropy is strictly positive and thus yields an IG metric that does not share this problem.
@@
Vendi scoring.  The VIG criterion is based on the Vendi Score (Friedman and Dieng, 2023; Pasarkar et al., 2023). Since its inception, the VS has been applied to various settings such as enhancing the diversity of generative models’ output (Berns et al., 2023), tuning large language models (Wu et al., 2023), encouraging exploration in molecular simulations (Pasarkar et al., 2023), evaluating cultural diversity of text-to-image models beyond existing criteria such as faithfulness and aesthetics (Hall et al., 2024; Senthilkumar et al., 2024), feature ranking (Mousavi and Khalili, 2024), evaluating diversity of prompt-based generative models (Jalali et al., 2024), developing novel RAG pipelines for LLMs (Rezaei and Dieng, 2025b), improving the robustness of sequence generative models to varying noise levels (Rezaei and Dieng, 2025a), and in Bayesian optimization (Liu et al., 2024; Nguyen and Dieng, 2024). Most recently, Pasarkar and Dieng (2025) proposed the Vendiscope, a method based on the VS to explore and analyze the composition of large-scale datasets. Here, we leverage the VS to propose a novel IG measure that addresses many of MI’s drawbacks and yields novel frameworks for active data acquisition and level-set estimation.
@@ 4 Discussion
5 Theory and methods
We present several properties of VIG, and subsequently the mathematical models in the analyses and the implementation of VIG in the respective applications.
5.1 Understanding the Vendi Information Gain
$K_{\bm{p}}=\mathrm{diag}(\bm{p}).$ (5)
[…]
$H_{V}(D_{x,y};q)=H_{V}(D_{x};q)+H_{V}(D_{y};q),$ (9)
$K_{x,y}=K_{x}\otimes K_{y}.$ (10)
and thus the Vendi entropy of $D_{x,y}$ may be computed as:
5.2 Visualizing mutual information’s failures
@@ 5.3 Communication channels
5.4 The human response time model
@@
In section 1, we perform the simulation with the model by Christie et al. (2023) by transmitting three symbols $A$, $B$, and $C$ with varying degrees of similarity. In particular, we construct the kernel matrix:
@@
5.5 The active data acquisition framework
@@
Code Availability
Data Availability
Acknowledgements
Appendix
@@ Acknowledgements
First, fig. 7 shows the distribution of response times by VIG under different orders $q$, where larger values of $q$ lead to shorter response times. This trend is explained by the effects of $q$ mentioned above: as small orders are more sensitive to rarity, more data, and thus more time, is required for IG about the most common class to exceed the threshold. figs. 8 and 9 show the results of VIG across a wide range of values for $q$ under active data acquisition tasks. We observe the reassuring trend that different versions of VIG behave similarly, demonstrating the VIG’s robustness against this hyperparameter.

### v2

@@ 1 Introduction
In Section 4, we demonstrate that VIG overcomes MI’s limitations and outperforms it in diverse settings. We also introduce in Section 4 a novel VIG-based active data acquisition framework, delivering the first fully information-theoretic method for level-set estimation. Finally, we derive VIG’s key properties in Section A.
2 Background
@@
3 Method
We now define VIG and present its properties as an IG metric.
3.1 Defining the Vendi Information Gain
@@
3.2 Understanding the Vendi Information Gain
$K_{\bm{p}}=\mathrm{diag}(\bm{p}).$ (5)
[…]
$H_{V}(D_{x,y};q)=H_{V}(D_{x};q)+H_{V}(D_{y};q),$ (9)
$K_{x,y}=K_{x}\otimes K_{y}.$ (10)
and thus the Vendi entropy of $D_{x,y}$ may be computed as:
4 Experiments
We now demonstrate VIG’s superiority over MI in several applications. In particular, we use VIG to develop a unified framework for active data acquisition, including a novel, first-ever information-theoretic approach to level-set estimation.
@@ 4.1 Modeling human response times to external stimuli
Figure 2: Simulated response times as a function of similarity between possible messages by MI and VIG against observed experimental data from Slamecka (1963). MI’s insensitivity to inter-message similarity prevents the simulation of different behaviors under different conditions, while VIG’s simulations closely match observed real-world data.
@@
We demonstrate VIG’s ability to model the concept of confusability by performing the simulation with the model by Christie et al. (2023), transmitting three symbols $A$, $B$, and $C$ with varying degrees of similarity. In particular, symbol $C$ is modeled as being distinct from $A$ and $B$, while the similarity between $A$ and $B$ varies from low to intermediate to high. We repeat the same simulation 100 times and show in fig. 2 the distribution of simulated response times resulting from using MI and VIG, as well as the real-world data reported by Slamecka (1963) from their word-meaning experiments as a reference. From the real-world experimental data, we observe the natural trend that the higher the similarity between the stimuli, the harder it is to distinguish between them, and thus the longer the response time is. This trend cannot be replicated using MI but is faithfully simulated by VIG, demonstrating VIG’s advantage.
@@ 4.2 Vendi information gain leads to a general unified framework for active data acquisition
Finally, we perform ablation studies on the effect of the order $q$ has on VIG in figs. 9 and 10 and find that performance remains fairly robust across a wide range of values for $q$, suggesting that VIG’s main advantage over other policies stems from its IG reasoning for active data acquisition tasks.
@@ 5 Related Work
Mutual information.  Previous works have investigated the failure modes of MI. Indeed, researchers have highlighted the challenge in accurately estimating MI in an unbiased way, especially in high dimensions (Macke et al., 2011; Dorval, 2011; Rhee et al., 2012; Kinney and Atwal, 2014; Cheng and Vlachos, 2024). Many studies have also found MI estimators to take on undesirable pathological behaviors under multimodal, non-Gaussian distributions (Czyż et al., 2023; Peyrard and Cho, 2025), and even the identification of these undesirable behaviors of MI is still an active area of search. Approximating MI may require exponentially many samples (Quian Quiroga and Panzeri, 2009), and even when it can be approximated with high fidelity or computed exactly, MI’s insensitivity to random variables’ spread may mischaracterize IG (Petty, 2018; Oliver, 2022; Cheng and Vlachos, 2024). We echo the findings by Schroeder (2004), who argued that MI is a good IG metric for communication systems when there is no need to model similarity between transmitted symbols, but fails in other contexts.
@@
Various works have considered alternatives to MI that address its drawbacks. However, these proposals are tailored to specific applications thus do not generalize to a variety of tasks where MI is employed. For instance, Torkkola (2003) considered a version of MI that accounts for the similarity between pairs of samples (something VIG inherently does) using a Parzen estimator as part of a feature extraction procedure. For image registration, Haber and Modersitzki (2005) calculated a distance-aware MI-based quantity and showed that this alternative outperforms MI under multi-modal data. In hierarchical clustering where MI has been found to inappropriately favor large clusters under high dimensions, Marrelec et al. (2015) used the empirical covariance matrix, like VIG, to correct the cluster-favoring behavior, while Hafemann et al. (2024) addressed MI’s insensitivity to spread by incorporating each data point’s distances to its nearest neighbors when analyzing LiDAR sensors. Meister et al. (2024) computed a similarity-weighted version of Shannon entropy to model contextual word predictability. Similar to findings in our response time application, modeling similarity between words better bridges the gap between information theoretic simulations and real-world data from cognitive science compared to MI. Liu et al. (2022) employed a distance-aware measure that also uses the covariance matrix of the samples to sidestep MI’s intractability for learning fair representations. Kala (2021) pointed out that differential entropy may take on negative values, resulting in unintuitive IG values for sensitivity analysis tasks under MI, and Petty (2018) made similar observations in the context of modeling atmospheric variables. Vendi entropy is strictly positive, and thus VIG does not share this problem.
@@
Vendi scoring.  The VIG criterion is based on the Vendi Score (Friedman and Dieng, 2023; Pasarkar et al., 2023). Since its inception, the VS has been applied to various settings such as enhancing the diversity of generative models’ output (Berns et al., 2023), tuning large language models (Wu et al., 2023), encouraging exploration in molecular simulations (Pasarkar et al., 2023), evaluating cultural diversity of text-to-image models beyond existing criteria such as faithfulness and aesthetics (Hall et al., 2024; Senthilkumar et al., 2024), feature ranking (Mousavi and Khalili, 2024), evaluating diversity of prompt-based generative models (Jalali et al., 2024), developing novel RAG pipelines for LLMs (Rezaei and Dieng, 2025b), improving the robustness of sequence generative models to varying noise levels (Rezaei and Dieng, 2025a), and Bayesian optimization (Liu et al., 2024; Nguyen and Dieng, 2024). Most recently, Pasarkar and Dieng (2025) proposed the Vendiscope, a method based on the VS to explore and analyze the composition of large-scale datasets. We leverage the VS to propose a novel IG measure that addresses many of MI’s drawbacks and yields novel frameworks for active data acquisition and level-set estimation.
@@ 6 Discussion
7 Code Availability
8 Data Availability
Acknowledgements
Appendix A Details on experiments
We present the mathematical models in the analyses and the implementation of VIG in the respective applications.
A.1 Visualizing mutual information’s failures
@@ A.2 Communication channels
A.3 The human response time model
@@
In fig. 7, we perform the simulation with the model by Christie et al. (2023) by transmitting three symbols $A$, $B$, and $C$ with varying degrees of similarity. In particular, we construct the kernel matrix:
@@
Figure 7: Progression of posterior probabilities and information gain, as measured by MI and VIG, under varying degrees of inter-message similarity as a function of time. MI plots the same trend across different conditions, while VIG becomes lower under more similar messages, appropriately quantifying increasing difficulty in distinguishing similar messages.
Symbol $A$ is set to be the intended stimulus, and the top panel of fig. 7 shows the posterior probability that each symbol is the intended stimulus $p(m\mid D_{t})$ as the function of time $t$. Here, we start out with a flat prior where each symbol has a $1/3$ probability of being the intended stimulus, and as we observe more data as time goes on, $p(m=A\mid D_{t})$ increases, indicating that $m=A$ is the true message.
The remaining panels of fig. 7 show the progression of IG as measured by MI and VIG in each of the three aforementioned settings. MI treats individual stimuli as completely distinct from one another and plots the same trend across the settings, failing to distinguish the different conditions. Meanwhile, VIG appropriately quantifies increasing difficulty in distinguishing messages as they become more similar, as demonstrated by lower IG. (In the extreme case of low similarity where there is no inter-stimulus similarity, VIG reduces to MI.)
A.4 The active data acquisition framework
@@
Appendix B Additional results
@@
First, fig. 8 shows the distribution of response times by VIG under different orders $q$, where larger values of $q$ lead to shorter response times. This trend is explained by the effects of $q$ mentioned above: as small orders are more sensitive to rarity, more data, and thus more time, is required for IG about the most common class to exceed the threshold. figs. 9 and 10 show the results of VIG across a wide range of values for $q$ under active data acquisition tasks. We observe the reassuring trend that different versions of VIG behave similarly, demonstrating the VIG’s robustness against this hyperparameter.